# Directional Textual Inversion for Personalized Text-to-Image Generation

### v1

@@ 4.4 Human evaluation
To further examine the effectiveness of our method, we conducted a large scale user study (100 participants via Amazon Mechanical Turk) to measure real-world user preferences. Each participant was asked to respond to 20 questions, comprising 10 questions assessing subject fidelity and 10 questions evaluating image-text alignment. Participants were instructed to select the output that best met the specified criteria for each question. To ensure the reliability of the study, we excluded four user responses that did not adhere to the specified instructions. A fixed random seed was employed, and the answer options were shuffled for each question. The results, summarized in Table 4, show that DTI consistently outperforms the other methods on both metrics, indicating that its improvements in alignment are clearly perceived by human evaluators. More details of this user study can be found in Appendix D.5.
Figure 4: We compare images generated by a TI and our DTI. Two personalized subjects are interpolated, including interpolation between inanimate and animate subjects, live subjects, and human faces. Images are generated with interpolation ratio [0.0,0.35,0.40,0.45,0.50,0.55,0.60,0.65,1.0][0.0,0.35,0.40,0.45,0.50,0.55,0.60,0.65,1.0] for better visualization. Our DTI offers smooth interpolation between concepts, expanding the personalization in more creative axis.
@@ 4.5 Embedding interpolation for creative applications
The advantages of our approach are clearly visible across different domains. As shown in the first rows of the figure, one can seamlessly merge a dog and a teapot, resulting in imaginative hybrid objects like an adorable teapot that progressively adopts the features of the dog. This indicates that DTI excels at blending conceptually distinct subjects, a significant creative application. In the second example, it can create the creative animal between a dog and a cat, that merges the features of each animal in a smooth manner. Lastly, DTI smoothly interpolates between the faces of a young boy and an older woman, generating a plausible progression that simultaneously alters age and appearance while maintaining facial coherence. This highlights its potential for nuanced face personalization.
@@ 5.1 Personalized text-to-image generation
Textual Inversion (gal_image_2023), which is most well-known for its lightweight integration to many other personalization works, uses embedding optimization by introducing learnable tokens for personalized information without model modification. Subsequent work explored diverse embedding strategies (voynov_p_2023; alaluf2023neural; wu2025core; zhang2024compositional), often with demanding excessive computational costs. Among them, CrossInit (pang2024cross) offered an efficient initialization strategy with minimal overhead, replacing initialization tokens with the output of text encoder and using regularization loss.
In contrast, fine-tuning based methods such as DreamBooth (ruiz_dreambooth_2023) achieve high subject fidelity, but require significant computational resources compared to embedding optimization methods (kumari_multi-concept_2023; han2023svdiff; gu2023mix; chen2023disenbooth; tewel2023key; zhang2024attention; qiu2023controlling; pang2024attndreambooth) . More recently, park2024textboost proposed fine-tuning text encoder instead of image generator for efficiency, but they still demand more parameters compared to embedding optimization methods.
@@ 5.2 Directional embedding space
A number of prior works has emphasized constraining embedding representations to the hypersphere. These include using vMF mixtures for directional clustering (jameel2019word), normalizing norms for face recognition (meng2019spherical), angle-optimized embeddings to address cosine saturation (li2024aoe), and spherical constraints for uniform document clustering (zhang2020deep). wang2020understanding offered theoretical support for hyperspherical constraints in contrastive learning. Our method aligns with this trend by modeling embeddings as directional distributions but uniquely decomposes and explicitly optimizes textual embedding direction using a vMF prior within Textual Inversion framework.
@@ 6 Discussion & Conclusion
Overall, our work tackles a key challenge in personalized text-to-image generation: achieving a strong alignment between text prompts and generated imagery. We have identified and rigorously analyzed embedding norm inflation as a significant bottleneck to this alignment, providing both theoretical and empirical evidence of its detrimental effects. In addition, our investigation focuses on the directional characteristics of the token embedding space, an area that has been comparatively underexplored in the literature, particularly when contrasted with the extensive research dedicated to the output embedding space of text encoders. Leveraging this key insight into the semantic significance of token embedding directionality, we proposed Directional Textual Inversion (DTI), a novel framework that keeps the embedding norm to in-distribution scale and solely optimizes the direction. We further reformulate the conventional Textual Inversion optimization process by incorporating directional priors. Our DTI demonstrably enhances prompt fidelity, thereby substantially improving the practicality of token embedding-based personalization and enabling innovative creative applications such as the smooth interpolation of learned concepts. We truly hope our work paves the way for more effective and versatile token embedding-based personalization within generative AI, unlocking enhanced capabilities for users to articulate their unique creative visions with greater precision and control.
@@
To ensure the full reproducibility of our research, we provide our complete source code, experimental details, and dataset information in the supplementary material, which will be made publicly available on GitHub upon publication. We utilized publicly available datasets, mostly from DreamBooth, FFHQ and StyleDrop, and our repository will include scripts for any necessary preprocessing. Also, all of the packages are explicitly stated in the pyproject.toml file of our code. All experiments were conducted on a single NVIDIA A6000 GPU, with a training per subject for approximately 7 minutes with SDXL-base and 30 minutes with SANA1.5-1.6B. All hyperparameters are explicitly defined in the Appendix, and also in the run files of our code to ensure transparency and ease of use.
@@ Appendix A Embedding norm and direction
study
studies, studying, research, bookclub,
reading, studied, sketches, measurements, thumbnail
U+3160, texanscheer, asober, instaweatherpro,
mydayin, premiosmtvmiaw, tairp, thepersonalnetwork, U+2412
writing
writer, write, written, writ,
writers, writings, recording, blogging, wrote
phdlife, poetryday, tomorrowspaper, urstrulymahesh,
@___, twitterkurds, asober, fakespeare, jamiedor
@@ B.2 Positional attenuation
so $\|\mathrm{LN}(m{\bm{v}}+{\bm{p}})-\mathrm{LN}(m{\bm{v}})\|=\sqrt{d}\,\frac{({\bm{I}}_{d}-{\bm{u}}{\bm{u}}^{\top}){\bm{C}}{\bm{p}}}{m\left\lVert{\bm{C}}{\bm{v}}\right\rVert}+\mathcal{O}(m^{-2})$, which is $\mathcal{O}(\left\lVert{\bm{p}}\right\rVert/m)$. ∎
@@ B.3 Residual stagnation
$\frac{\|{\bm{x}}^{(\ell+1)}-{\bm{x}}^{(\ell)}\|_{2}}{\|{\bm{x}}^{(\ell)}\|_{2}}\leq\frac{B_{\ell}}{\|{\bm{x}}^{(\ell)}\|_{2}},\qquad\angle({{\bm{x}}^{(\ell)}},{{\bm{x}}^{(\ell+1)}})\leq\arcsin\!\Bigl(\frac{B_{\ell}}{\|{\bm{x}}^{(\ell)}\|_{2}}\Bigr).$
@@
Let $\theta_{\ell}:=\angle(x^{(\ell)},x^{(\ell+1)})$. By the recall above, $\theta_{\ell}\leq\arcsin\!\big(B_{\ell}/\left\lVert x^{(\ell)}\right\rVert\big)\leq\tfrac{\pi}{2}\,B_{\ell}/\left\lVert x^{(\ell)}\right\rVert$. Also $\left\lVert x^{(\ell)}\right\rVert\geq\left\lVert x^{(0)}\right\rVert-\sum_{j<\ell}B_{j}$ (each step can shrink the norm by at most $B_{\ell}$). Summing angles (spherical triangle inequality) gives the first display; since $\left\lVert x^{(0)}\right\rVert-\sum_{j<\ell}B_{j}\geq\left\lVert x^{(0)}\right\rVert-S_{L}$, each fraction is $\leq B_{\ell}/(\left\lVert x^{(0)}\right\rVert-S_{L})$, yielding the last bound. ∎
@@ C.2 Why vMF over other distributions?
We chose the von Mises-Fisher (vMF) distribution as it is ideally suited for modeling the directional characteristics of token embeddings we identified in Section 2. Our central hypothesis is that the token embedding vocabulary can be modeled as a mixture of vMF distributions, where each component corresponds to a distinct semantic cluster (e.g., one for animals, another for objects). The vMF distribution is the suitable building block for this model for three key reasons:
@@ D.1 Implementation Details
Following the protocol of recent studies, we primarily conducted experiments using Stable Diffusion XL (SDXL). To demonstrate broader applicability to different models, we also conducted experiments with very recent model, SANA 1.5 (xie2024sana), where the results can be found in Table 3.
For a fair comparison, we adopted most of the hyperparameter settings from the Textual Inversion (TI) implementation provided by the HuggingFace diffusers library. Specifically, we used a training batch size of 4, and enabled mixed-precision training with the bfloat16 (bf16) format. We set the learning rate commonly-used 5e35e-3. All experiments were run with a fixed random seed of 42, and the maximum number of training steps was set to 500. For output generation, we used the DDIMScheduler with 50 inference steps for SDXL and 20 steps with FlowMatchEulerDiscreteScheduler for SANA.
Hyperparemeters.  There can be various approaches to selecting the concentration parameter $\kappa$. We performed a grid search and found that values in the range of 5e-5 to 2e-4 works well. Therefore, we did not conduct a extensive search for an optimal decimal value. Throughout the experiments, we simply fixed value to 1e-4, which generalizes well to experiments with different settings. Examples illustrating the effects of different $\kappa$ settings are provided in Table 3.
Baselines.  Throughout this paper, we compare our method with two baseline approaches: Textual Inversion (TI) (gal_image_2023) and CrossInit (pang2024cross). Since the official CrossInit implementation is based on Stable Diffusion v2.1 with hyperparameters tailored to that version, we reconfigure it to operate on SDXL by aligning its training setup with that of TI. Specifically, we adopt the same hyperparameters as used for TI, and we set the regularization weight for CrossInit to 1e51e-5, as specified in the original paper.
@@ D.2 On the choice of prior
For all of our experiments in the main section, we used the initial tokens as prior from the DreamBooth dataset as is. However, we would like to note that since our DTI can leverage the prior, searching for better priors can lead to better results. This demonstrates the effectiveness of the prior.
@@ D.3 Comparison with other baselines
Although DTI is primarily designed for embedding-only personalization, it also functions effectively as a drop-in replacement within fine-tuning pipelines. Recent work on Direct Consistency Optimization (DCO) lee_direct_2024 typically performs joint optimization of a concept token using standard Textual Inversion (TI). To assess the impact of substituting this TI component with DTI, we conducted joint training for 250 steps using a LoRA module with rank 4.
@@
Figure 6: Qualitative results with DCO (lee_direct_2024). We provide the comparison of the output images of DCO + TI and DCO + DTI. The results suggest our DTI’s superiority in text fidelity while preserving strong image similarity.
@@ D.4 Ablation study
  • Setting $\kappa=0.0$, which removes the directional prior, leads to lower scores in text fidelity, which validates our method’s priority in model’s enhancing semantic understanding.
@@
Overall, our default choice of $\kappa=1e-4$ provides a better balance between maintaining subject similarity and ensuring text fidelity. Note that $\kappa=1e-4$ may not be strictly optimal in decimals across all criteria but works reasonably well by providing robust overall performance.
@@ D.6 More qualitative results
Figure 8: Qualitative results with SDXL. Here, we provide more qualitative comparison with TI and CrossInit. Our DTI consistently generates results that precisely reflect the user text prompts, maintaining the subject similarity at the same time.
Figure 9: Qualitative results with SANA1.5-1.6B. Here, we provide more qualitative comparison with TI and CrossInit on SANA. Our DTI consistently generates results that precisely reflect the user text prompts, maintaining the subject similarity at the same time.
@@
Figure 11: Qualitative results with TI/DTI with LoRA on SDXL. We have performed qualitative comparison of applying TI and DTI on model fine-tuning methods using LoRA (rank 32). DTI consistently improves the text prompt fidelity compared to TI.
@@ D.7 More results on applications
My object in my style.  We also compare our results in simultaneous generation of personalized subject and style. The results demonstrated in Figure 13 shows that DTI successfully generates outputs that are faithful to both subject and style, while TI fails to.
@@ Appendix E Additional Experiments
Figure 16: Ablation on magnitude settings. For consistency and ease of use, all magnitudes in this paper are set to the average value computed over the model’s vocabulary (see ablation in Table 3). To evaluate the effect of using concept-specific magnitudes (e.g., the magnitude of ‘cat’ for the concept ¡cat¿), we provide ablation results under different magnitude settings. The results show that small deviations from the default magnitude do not lead to noticeable differences in output quality.

### v2

@@ 4.4 Human evaluation
To further examine the effectiveness of our method, we conducted a large-scale user study (100 participants via Amazon Mechanical Turk) to measure real-world user preferences. Each participant was asked to respond to 20 questions, comprising 10 questions assessing subject fidelity and 10 questions evaluating image-text alignment. Participants were instructed to select the output that best met the specified criteria for each question. To ensure the reliability of the study, we excluded four user responses that did not adhere to the specified instructions. A fixed random seed was employed, and the answer options were shuffled for each question. The results, summarized in Table 4, show that DTI consistently outperforms the other methods on both metrics, indicating that its improvements in alignment are clearly perceived by human evaluators. More details of this user study can be found in Appendix D.5.
Figure 4: We compare images generated by a TI and our DTI. Two personalized subjects are interpolated, including interpolation between inanimate and animate subjects, live subjects, and human faces. Images are generated with interpolation ratio [0.0,0.35,0.40,0.45,0.50,0.55,0.60,0.65,1.0][0.0,0.35,0.40,0.45,0.50,0.55,0.60,0.65,1.0] for better visualization. Our DTI offers smooth interpolation between concepts, expanding personalization along a more creative axis.
@@ 4.5 Embedding interpolation for creative applications
The advantages of our approach are clearly visible across different domains. As shown in the first rows of the figure, one can seamlessly merge a dog and a teapot, resulting in imaginative hybrid objects like an adorable teapot that progressively adopts the features of the dog. This indicates that DTI excels at blending conceptually distinct subjects, a significant creative application. In the second example, it can create a creative animal between a dog and a cat, that merges the features of each animal in a smooth manner. Lastly, DTI smoothly interpolates between the faces of a young boy and an older woman, generating a plausible progression that simultaneously alters age and appearance while maintaining facial coherence. This highlights its potential for nuanced face personalization.
@@ 5.1 Personalized text-to-image generation
Textual Inversion (gal_image_2023), which is most well-known for its lightweight integration into many other personalization works, uses embedding optimization by introducing learnable tokens for personalized information without model modification. Subsequent work explored diverse embedding strategies (voynov_p_2023; alaluf2023neural; wu2025core; zhang2024compositional), often at excessive computational cost. Among them, CrossInit (pang2024cross) offered an efficient initialization strategy with minimal overhead, replacing initialization tokens with the output of text encoder and using a regularization loss.
In contrast, fine-tuning based methods such as DreamBooth (ruiz_dreambooth_2023) achieve high subject fidelity, but require significant computational resources compared to embedding optimization methods (kumari_multi-concept_2023; han2023svdiff; gu2023mix; chen2023disenbooth; tewel2023key; zhang2024attention; qiu2023controlling; pang2024attndreambooth) . More recently, park2024textboost proposed fine-tuning the text encoder instead of the image generator for efficiency, but they still demand more parameters compared to embedding optimization methods.
@@ 5.2 Directional embedding space
A number of prior works have emphasized constraining embedding representations to the hypersphere. These include using vMF mixtures for directional clustering (jameel2019word), normalizing norms for face recognition (meng2019spherical), angle-optimized embeddings to address cosine saturation (li2024aoe), and spherical constraints for uniform document clustering (zhang2020deep). wang2020understanding offered theoretical support for hyperspherical constraints in contrastive learning. Our method aligns with this trend by modeling embeddings as directional distributions but uniquely decomposes and explicitly optimizes textual embedding direction using a vMF prior within the Textual Inversion framework.
@@ 6 Discussion & Conclusion
Overall, our work tackles a key challenge in personalized text-to-image generation: achieving a strong alignment between text prompts and generated imagery. We have identified and rigorously analyzed embedding norm inflation as a significant bottleneck to this alignment, providing both theoretical and empirical evidence of its detrimental effects. In addition, our investigation focuses on the directional characteristics of the token embedding space, an area that has been comparatively underexplored in the literature, particularly when contrasted with the extensive research dedicated to the output embedding space of text encoders. Leveraging this key insight into the semantic significance of token embedding directionality, we proposed Directional Textual Inversion (DTI), a novel framework that keeps the embedding norm at an in-distribution scale and solely optimizes the direction. We further reformulate the conventional Textual Inversion optimization process by incorporating directional priors. Our DTI demonstrably enhances prompt fidelity, thereby substantially improving the practicality of token embedding-based personalization and enabling innovative creative applications such as the smooth interpolation of learned concepts. We truly hope our work paves the way for more effective and versatile token embedding-based personalization within generative AI, unlocking enhanced capabilities for users to articulate their unique creative visions with greater precision and control.
Acknowledgments
This research was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (No. RS-2025-00520207); the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (Nos. 2022-0-00680, 2022-0-01045, RS-2024-00457882, RS-2025-02217259, RS-2019-II190075), including the National AI Research Lab Project and the Artificial Intelligence Graduate School Support Program (KAIST); and the Korea Evaluation Institute of Industrial Technology (KEIT) grants funded by the Korea government (MOTIE) (Nos. 2022-0-00680, 2022-0-01045, RS-2025-02217259).
@@ Acknowledgments
To ensure the full reproducibility of our research, we provide our complete source code, experimental details, and dataset information. We utilized publicly available datasets, primarily from DreamBooth (ruiz_dreambooth_2023), FFHQ (karras2019style), and StyleDrop (sohn2023styledrop), and our repository includes scripts for all necessary preprocessing. Additionally, all software dependencies are explicitly specified in the pyproject.toml file. All experiments were conducted on a single NVIDIA A6000 GPU, with training taking approximately 7 minutes per subject for SDXL-base and 30 minutes for SANA 1.5-1.6B. To guarantee transparency and ease of use, all hyperparameters are detailed in the Appendix and are also included in the run scripts within our codebase.
@@ Appendix A Embedding norm and direction
study studies, studying, research, bookclub, reading, studied, sketches, measurements, thumbnail U+3160, texanscheer, asober, instaweatherpro, mydayin, premiosmtvmiaw, tairp, thepersonalnetwork, U+2412
writing writer, write, written, writ, writers, writings, recording, blogging, wrote phdlife, poetryday, tomorrowspaper, urstrulymahesh, @___, twitterkurds, asober, fakespeare, jamiedor
@@ B.2 Positional attenuation
so $\|\mathrm{LN}(m{\bm{v}}+{\bm{p}})-\mathrm{LN}(m{\bm{v}})\|\leq\sqrt{d}\,\frac{\|({\bm{I}}_{d}-{\bm{u}}{\bm{u}}^{\top}){\bm{C}}{\bm{p}}\|}{m\left\lVert{\bm{C}}{\bm{v}}\right\rVert}+\mathcal{O}(m^{-2})$, which is $\mathcal{O}(\left\lVert{\bm{p}}\right\rVert/m)$. ∎
@@ B.3 Residual stagnation
$\frac{\|{\bm{x}}^{(\ell+1)}-{\bm{x}}^{(\ell)}\|_{2}}{\|{\bm{x}}^{(\ell)}\|_{2}}\leq\frac{B_{\ell}}{\|{\bm{x}}^{(\ell)}\|_{2}}.$
If additionally $B_{\ell}<\|{\bm{x}}^{(\ell)}\|_{2}$,
$\angle({{\bm{x}}^{(\ell)}},{{\bm{x}}^{(\ell+1)}})\leq\arcsin\!\Bigl(\frac{B_{\ell}}{\|{\bm{x}}^{(\ell)}\|_{2}}\Bigr).$
@@
Let $\theta_{\ell}:=\angle(x^{(\ell)},x^{(\ell+1)})$. By Lemma 2 above, $\theta_{\ell}\leq\arcsin\!\big(B_{\ell}/\left\lVert x^{(\ell)}\right\rVert\big)\leq\tfrac{\pi}{2}\,B_{\ell}/\left\lVert x^{(\ell)}\right\rVert$. Also $\left\lVert x^{(\ell)}\right\rVert\geq\left\lVert x^{(0)}\right\rVert-\sum_{j<\ell}B_{j}$ (each step $j$ can shrink the norm by at most $B_{j}$). Summing angles (spherical triangle inequality) gives the first display; since $\left\lVert x^{(0)}\right\rVert-\sum_{j<\ell}B_{j}\geq\left\lVert x^{(0)}\right\rVert-S_{L}$, each fraction is $\leq B_{\ell}/(\left\lVert x^{(0)}\right\rVert-S_{L})$, yielding the last bound. ∎
@@ C.2 Why vMF over other distributions?
We chose the von Mises-Fisher (vMF) distribution as it is ideally suited for modeling the directional characteristics of token embeddings we identified in Section 2. Our central hypothesis is that the token embedding vocabulary can be modeled as a mixture of vMF distributions, where each component corresponds to a distinct semantic cluster (e.g., one for animals, another for objects). The vMF distribution is a suitable building block for this model for three key reasons:
@@ D.1 Implementation Details
Following the protocol of recent studies, we primarily conducted experiments using Stable Diffusion XL (SDXL). To demonstrate broader applicability to different models, we also conducted experiments with the recent model SANA 1.5 (xie2025sana); these results are presented in Table 3.
For a fair comparison, we adopted most hyperparameter settings from the Textual Inversion (TI) implementation provided by the Hugging Face diffusers library (von-platen-etal-2022-diffusers). We used a training batch size of 4 and enabled mixed-precision training with the bfloat16 (bf16) format. For the baselines, the learning rate was set to the commonly used value of $5\times 10^{-3}$. Since our method employs a different optimizer, we conducted a separate learning rate search and set it to $0.02$. Following rombach_high-resolution_2022 and gal_image_2023, we scaled the base learning rate proportionally to the batch size for all methods.
All experiments were performed with a fixed random seed of 42, and the maximum number of training steps was set to 500 for SDXL and 1000 for SANA. For image generation, we used the DDIMScheduler with 50 inference steps for SDXL and the FlowMatchEulerDiscreteScheduler with 20 inference steps for SANA.
Hyperparameters.  There are various approaches to selecting the concentration parameter $\kappa$. We performed a grid search and found that values in the range of $5\times 10^{-5}$ to $2\times 10^{-4}$ work well. Therefore, we did not conduct an exhaustive search for a more precise optimal value. Throughout our experiments, we simply fixed $\kappa$ at $10^{-4}$, which generalizes well across different settings. Examples illustrating the effects of varying $\kappa$ are provided in Table 3.
Baselines.  Throughout this paper, we compare our method with two baseline approaches: Textual Inversion (TI) (gal_image_2023) and CrossInit (pang2024cross). Since the official CrossInit implementation is based on Stable Diffusion v2.1 with hyperparameters tailored to that version, we reconfigure it to operate on SDXL by aligning its training setup with that of TI. Specifically, we adopt the same hyperparameters as used for TI, and we set the regularization weight for CrossInit to $10^{-5}$, as specified in the original paper.
@@ D.2 On the choice of prior
For all of our experiments in the main section, we used the same class tokens as those specified in the DreamBooth dataset as priors. However, we would like to note that since our DTI can leverage the prior, searching for better priors can lead to better results. This demonstrates the effectiveness of the prior.
@@ D.3 Comparison with other baselines
Although DTI is primarily designed for embedding-only personalization, it also functions effectively as a drop-in replacement within fine-tuning pipelines. Recent work on Direct Consistency Optimization (DCO; lee_direct_2024) typically performs joint optimization of a concept token using standard Textual Inversion (TI). To assess the impact of substituting this TI component with DTI, we conducted joint training for 250 steps using a LoRA module with rank 4.
@@
Figure 6: Qualitative results with DCO (lee_direct_2024). We provide a comparison of the output images of DCO + TI and DCO + DTI. The results suggest our DTI’s superiority in text fidelity while preserving strong image similarity.
@@ D.4 Ablation study
  • Setting $\kappa=0.0$, which removes the directional prior, leads to lower scores in text fidelity, which suggests that the directional prior improves semantic alignment.
@@
Overall, our default choice of $\kappa=1e-4$ provides a better balance between maintaining subject similarity and ensuring text fidelity. Note that $\kappa=1e-4$ may not be numerically optimal across all criteria but works reasonably well by providing robust overall performance.
@@ D.6 More qualitative results
Figure 8: Qualitative results with SDXL. Here, we provide more qualitative comparisons with TI and CrossInit. Our DTI consistently generates results that precisely reflect the user text prompts, while maintaining subject similarity.
Figure 9: Qualitative results with SANA1.5-1.6B. Here, we provide more qualitative comparisons with TI and CrossInit on SANA. Our DTI consistently generates results that precisely reflect the user text prompts, maintaining subject similarity.
@@
Figure 11: Qualitative results with TI/DTI with LoRA on SDXL. We provide a qualitative comparison of TI and DTI when combined with LoRA-based fine-tuning (rank 32). DTI consistently improves the text prompt fidelity compared to TI.
@@ D.7 More results on applications
My object in my style.  We also compare our results in simultaneous generation of personalized subject and style. The results demonstrated in Figure 13 show that DTI successfully generates outputs that are faithful to both subject and style, while TI fails to.
@@ Appendix E Additional Experiments
Figure 16: Ablation on magnitude settings. For consistency and ease of use, all magnitudes in this paper are set to the average value computed over the model’s vocabulary (see ablation in Table 3). To evaluate the effect of using concept-specific magnitudes (e.g., the magnitude of ‘cat’ for the concept <cat>), we provide ablation results under different magnitude settings. The results show that small deviations from the default magnitude do not lead to noticeable differences in output quality.